- 208: Radiation, Site stn1, Fri 14-Dec-2001 11:35:04 MST, Tear-down documentation of radiometers
This information has been transcribed from my CASES97/99 field notebook.
At tear-down I went around the four surface radiation sites and removed
the radiometers and the soil sensors. I made field notes at the sites.
The notes were made on Saturday, 30 Oct''99, or possible the next day.
Station 1
pyrg.in 29260 net 97034 Htfluxplate 87036
pyrg.out 26416 SoilT 001
Soilq 001
Rainguage 13
Station 2
pyrg.in 31977 net 94197? Htfluxplate 943149
pyrg.out 31981 K&Zin 940181 SoilT NCAR 002
K&Zout 940187 Soilq 009
Rainguage 3
Station 3
pyrg.out 31975 net 94125 Htfluxplate H983212
SoilT NCAR 003
Soilq 6
Rainguage 32
Station 5
pyrg.out 31979 net 97035
- 202: Radiation, Site base, Thu 10-Feb-2000 11:25:28 MST, Pyrgeometer configuration in CASES99
The ten pyrgeoemeters used during CASES99 were set out in two configurations:
the intercomparison with the pigs mounted side-by side, all up-looking,
the measurement with the pigs dispersed to their specific location;
two up-looking at 50m on the tower boom,
two down-looking at 50 m on the tower boom,
two uplooking at 2m, one at station 1, one at station 2,
four down-looking at 2m, one at station 1, one at station 2,
one at station 3 and one at station 5.
The parameter identifications corresponding to these two configurations are:
Intercomparison Measurement
Easterly-most
26416 (Rpile,Tcase,Tdome).out.s02 (Rpile,Tcase,Tdome).out.s2 Bluestem
29260 (Rpile,Tcase,Tdome).in.s02 (Rpile,Tcase,Tdome).in.s2 Bluestem
31977 (Rpile.Tcase,Tdome).in.s01 (Rpile,Tcase,Tdome).in.s1 Shortgrass
31981 (Rpile,Tcase,Tdome).out.s01 (Rpile,Tcase,Tdome).out.s1 Shortgrass
31975 (Rpile,Tcase,Tdome).out.s05 (Rpile,Tcase,Tdome).out.s5 Weed
31974 (Rpile,Tcase,Tdome).out.a.boom (Rpile,Tcase,Tdome).out.a.50m Tower 50m
31980 (Rpile,Tcase,Tdome).in.a.boom (Rpile,Tcase,Tdome).in.a.50m Tower 50m
31976 (Rpile,Tcase,Tdome).in.b.boom (Rpile,Tcase,Tdome).in.b.50m Tower 50m
31978 (Rpile,Tcase,Tdome).out.b.boom (Rpile,Tcase,Tdome).out.b.50m Tower 50m
31979 (Rpile,Tcase,Tdome).out.s03 (Rpile,Tcase,Tdome).out.s3 Drainage
Westerly-most
The pyrgeometers were in intercomparison configuration:
1999 September 25, ??:00 to 1999 October 4, ??:00
The pyrgeometers were in measurement configuration :
1999 October 4, ??:00 to 1999 October 30, ??:00
For more details, including schematics, see:
/net/sssf1/delany/deployments/cases99/lwdiverg/lwdivanal.fm
- 201: Radiation, Site , Wed 08-Dec-1999 10:34:02 MST, Snow cover on pyrgs
At 1030 Wed 8 Dec''99 I temporarilly depowered the fans to allow a full
snow cover over all the pyrgs.
It started to snow in the nightime and although the fans had kept the dome
clear snow had built up on the shield so that the snow formed an annulus
higher than the top of the domes.
Snow is forecast to continue for the next 24 hours
The extra snow did not occur, the existing snow on the domes evaporated.
On 9 Dec at 11-30 I restarted the fans and cleaned the domes.
- 200: Radiation, Site , Fri 03-Dec-1999 15:57:28 MST, Roof-top post-CASES99 pyrg-cal
The boom pigs' data has been plagued by intermittent spikes
Today, at ~1030 I replaced the serial cable from the boom-box Campbel
to the serial BOB with a new one. This seems to have cured the problem.
At ~1700 I removed all the snow from the radiometers and cleaned the domes.
- 199: Radiation, Site , Wed 01-Dec-1999 09:15:32 MST, Pyrgeoemter Campbell coefficients
These are the coefficients determined from the oil-bath temp calibration of
the pyrgeoemters, prior to the CASES99 deployment. These coefficients were
used during the pre-CASES99 intercomparison in the field and during the
deployment. They are also to be used for the post-CASES99 intercomparison
on the roof at the Foothills Lab.
This logbook table was copied from
/net/sssf1/delany/deployments/cases99/lwdiverg/lbpigcoeff.txt
and then the non-pyrgeoemetrs coefficients were added
Serial#
Pile
CR10 coeff
Scaled Case coeff B0 x 10^0
Scaled Case coeff B1 x 10^3
Scaled Case coeff B2 x 10^6
Scaled Case coeff B3 x 10^9
Scaled Dome coeff B0 x 10^0
Scaled Dome coeff B1 x 10^3
Scaled Dome coeff B2 x 10^6
Scaled Dome coeff B3 x 10^9
Dome/ Case const
Connected to the Boom-Box Campbell
pyg.out.1
31974
3.8535
259.50435
.00335467
0.286452
-2.46374
3868.8
.00303966
.292687
-7.40681
2207.35
3.95
pyg.out.2
31978
3.5605
28031978
3.5605
280.85943
.00335437
0.287099
-2.0547
3531.01
.00303889
0.291274
-5.99237
1874.43
4.00
pyg.in.1
31980
3.6095
277.04668
.00335494
0.283676
1.33836
1980.11
.00305007
0.27288
4.55754
-137.92
3.90
pyg.in.2
31976
3.852
259.6054
.00335494
0.283467
1.8651
1587.36
.00304992
0.273924
3.66477
18.4609
3.80
Connected to the Sta 5 Campbell
pyg.out
31975
4.0275
248.29299
.00335599
.288404
-6.90295
6038.71
.00303124
0.30936
-16.7272
3864.95
3.05
Connected to the Sta 3 Campbell
pyg.out
31979
3.902
256.27883
.00335423
0.287392
-4.79718
4588.33
.00302993
0.307588
-15.2242
3443.34
3.20
Connected to the Sta 1 Campbell
pyg.in
31977
3.729
268.16841
.00335531
0.283141
2.81226
898.827
.00304622
0.280377
-1.0274
1105.48
3.50
pyg.out
31981
3.743
267.16538
.00335403
0.283362
2.6255
1028.04
.00304215
0.286235
-4.41592
1743.28
3.20
Connected to the Sta 2 Campbell
pyg.out
26416
(3.570)
280.11
.00335452
0.282295
2.92866
954.68
.00304556
0.282923
-3.22195
1552.84
(3.56)
pyg.in
29260
(3.210)
311.53
.00335383
0.27965
3.92473
597.175
.00306136
0.253163
15.4003
-2315.04
(3.56)
- 198: Radiation, Site , Tue 30-Nov-1999 15:45:48 MST, pyrgeometer coefficients
lwdiv
- 195: Radiation, Site base, Fri 19-Nov-1999 16:17:42 MST, Baud rates of rad data loggers
NOTE!
The campbell loggers on the PAM stations transmit at 1200 baud.
They were set this way in the field at CASES99 after the
intercomparison period.
- 186: Radiation, Site , Thu 28-Oct-1999 12:28:53 CDT, Cleaned radiometer domes
Between 1030 and 1100 the radiometer domes were cleaned at Stations 1, 2, 5 and 3. All the fans were operating smoothly
- 177: Radiation, Site , Tue 26-Oct-1999 20:02:28 CDT, Cleaned 50 m radiometer domes
At ~ 1530 on 26 October the 50 m trolley was pulled in and the four radiometers'
domes were cleaned with water wetted tissue. There was indication of brown dirt
on the domes.
- 174: Radiation, Site , Tue 26-Oct-1999 13:06:08 CDT, Cleaning domes
Between 1030 and 1115 the radiometer domes and the net radiometers were cleaned at Stations 1,3, 5 and 2. Photographs of the ground surfaces under the radiometers were taken.
- 169: Radiation, Site , Sun 24-Oct-1999 12:59:33 CDT, Cleaned radiometer domes
Between 1230 and 1300 the domes of the radiometers at stns, 1,2,5 and 3
were cleaned
- 161: Radiation, Site stn2, Fri 22-Oct-1999 13:33:22 CDT, Cleaning radiometers
At 11:15 the domes of the radiometers at station 1 were cleaned
At ~13:00 the domes of the radiometers at stations 2, 3, and 5 were cleaned.
There were traces of brownish grey dust, especially on station 1 domes
All net radiometers were cleaned.
All fans were running. The fan at station 3 was noisy.
- 136: Radiation, Site tower, Mon 18-Oct-1999 16:38:11 CDT, Cleaned radiometers at 50m and PAM stations
We retracted the 50m radiometers and cleaned them, beginning at 1500 CDT,
Oct 18. They were cleaned with "de-mineralized" water and kimwipes.
The radiometers were returned to their outboard position at 1615 CDT
(after unfouling the hauling line with the loooon green hook).
The domes of the surface radiation sensors were also cleaned after
finishing the tower radiometers.
- 135: Radiation, Site stn2, Mon 18-Oct-1999 11:59:54 CDT, PYG.OUT at station 2 questioned and fixed (?)
The pyg.out pile was bad so I went out to look at it.
According to Tom, this problem has happened before.
The voltage going to the multiplexer board was varying
quite a bit. I played with the connector inside the Campbell
which resulted in a good signal. The cables had been reversed
from a previous test so I switched them back to the correct
configuration. Will need to monitor to see if fix works. If
so, we will need to flag this Campbell box for repairs.
- 88: Radiation, Site stn2, Tue 12-Oct-1999 17:42:32 CDT, Rpile.out at Station 2
The Rpile.out data has been intermittent at station 2. At 3:50 pm,
local time, Oct 12, I disconnected and reconnected the cable at the
sensor and this appeared to solve the problem. At that time, I also
bumped the Schwartz out of level by popping the leveling screws out
of their upper seats. I reseated the screws and found that the level
was still okay. At that time I also adjusted the level of the net
radiometer about its N/S axis.
The Rpile.out data soon failed again and at 5:34 I swapped the pile.in
and pile.out cables at both ends. This fixed the problem again.
- 56: Radiation, Site stn1, Fri 08-Oct-1999 12:32:09 CDT, Corrections for radiometer deployment errors
I inventoried the surface radiometers to investigate the cause for
some anomolous long wave data:
Station Rnet Pyg.in Pyg.out Psp.in Psp.out
1 97034 29260 26416
2 94197 31977 31981 940181 940187
3 94125 31975
5 97035 31979
This showed that the radiometer arrays on stations 1 and 2 were
interchanged when deploying them at the end of the intercomparison
period. This morning (Oct 8) we entered into the CR10 loggers at
stations 1 and 2 the coefficients corresponding to the radiometers
mounted there.
WARNING WARNING WARNING WARNING WARNING WARNING
Be aware that the labeling of the radiometers at stations 1 and 2
is different for the intercomparison than it is after the radiometers
were deployed at the various PAM stations.
WARNING WARNING WARNING WARNING WARNING WARNING
The coefficients for the PSPs at station 2 were entered around 9:30
am local time.
The coefficients for the PYGs at station 1 were entered around 10:55
am local time.
The coefficients for the PYGs at station 2 were entered around 11:10
am local time. At that time we found that the signal cables for the
two PSPs were interchanged and this was corrected at the CR10.
At station 5 we found a small error in the third coefficient of the
Tcase.pyg.out calibration. The data sheet lists a coefficient of
-6.90295 and the CR10 has a coefficient of -6.9229. It was decided
to leave this coefficient as we found it, since this value was used
during the field intercomparison of the radiometers.
- 53: Radiation, Site , Fri 08-Oct-1999 08:52:04 CDT, Rnet cal factors
Rnet calibrations from REBS 8/17/99 calibration are:
Positive Negative (Zero Wind CF in W/m2 mV)
94196 9.25 11.44
97035 9.28 11.54 stn5 (stn5 is using these values)
94197 9.07 11.25 (stn2 is using these values)
91096 8.54 10.59
94125 8.65 10.75 (stn3 is using these values)
97034 9.31 11.39 stn1 (stn1 is using these values)
- 49: Radiation, Site , Wed 06-Oct-1999 17:07:04 CDT, radiometer work
Stuff done today:
stn1: changed CR10 baud rate to 1200. Commands are:
*0
7H
*1
A (until step 62)
51 (overwrites 52)
*0
(this can be done through EVE by typing:
entersys
stop
talk (port 4, rate 9600, etc.)
start
Rnet is Q97034
stn2: moved solar panels; hooked up battery swiped from stn5; mounted Rnet
with 2 screws (only had one before); leveled Rnet; added connector
to bare wires from Rnet so it could plug in; changed baud rate to 1200.
Rnet is labeled "station2"
stn3: moved solar panels; moved rad stand from IC array; leveled and oriented
stand; mounted Rnet from box labeled "station3", but with a label of
"station1"! Found that fan cable wires were color-coded backwards, so
rebuilt this cable (red negative, black positive); changed baud rate
to 1200.
stn5: stole battery from main station to power radiometers at stn2; removed
wrong DC power cable; attached good DC power cable; cut cable ties;
changed baud rate to 1200.
Rnet is Q97035.
All rad stand solar panels are now located about 4m to the NE of the CR10 box,
with the CR10 on the East side of the rad stand.
- 47: Radiation, Site , Tue 05-Oct-1999 19:38:10 CDT, moved rads out of IC
Installed 48m boom about 1300 today - leveled to within 0.003 deg. The
pull rope is tight! Also, don't trip over the CR10 box on the floor
at 49 m.
Mounted radiometers at stns 5, 2, and 1 (in that order), but not reporting yet.
stn5 is down waiting for a power cable/solar power move and baud rate change.
stn1 is waiting for a baud rate change
stn2 is waiting for a new battery!
stn3 is waiting for the whole rad stand to be moved.
- 41: Radiation, Site , Fri 01-Oct-1999 23:04:50 CDT, cleaned domes
Just cleaned domes on all pygs in the IC array. They all were somewhat
dirty. Also checked levels and all were good.
- 36: Radiation, Site stn3, Wed 29-Sep-1999 15:58:52 CDT, Tsfc replaced on PAM-3
Everest IRT Swapped:
Sensor 3577-4 removed; sensor 3577-2 installed
After the thunder/rain storm 2 nights ago the old sensor
was reading 100 deg-C.
- 33: Radiation, Site , Tue 28-Sep-1999 18:08:51 CDT, playing with rad biases from IC
Here is a first cut at adjusting biases on Rpile from the radiometers
based on data from the interval "99 Sep 25 17:00",lenhr=31,stns=0:
Rpile.in.a.boom Rpile.in.b.boom Rpile.in.s01 Rpile.in.s02 Rpile.out.a.boom
0 0.5 -1.2 bad 0.7
Rpile.out.b.boom Rpile.out.s01 Rpile.out.s02 Rpile.out.s03 Rpile.out.s05
0.4 -1 bad 0.2 10
- 30: Radiation, Site , Mon 27-Sep-1999 11:55:30 CDT, releveled radiometers
I just used the small bubble level to check the levels. Most looked okay,
though I adjusted the western 2 and the eastern 1 a bit.
- 25: Radiation, Site , Sat 25-Sep-1999 13:29:36 CDT, Pyg IC started!
The pyrgeometer intercomparison can be considered to have officially
started at 12:15 (1715 Z) today. Yesterday, Tony mounted fans on
all housings. Today, I just connected the radiometers which were
recalibrated (after their swim in oil) and received here last night.
(It was necessary to reposition both radiometers on their 8-hole base
plates to fit the housings.)
I also just leveled all of the Schwartz's (not knowing if this had
been done already) using the old digital level. Obviously, I didn't
have a good reference surface, but I still think that I got them
to within 0.1 degree. (They all sight to the horizon similarly.)
- 23: Radiation, Site , Fri 24-Sep-1999 10:59:27 MDT,
Replaced K&Z 970377 with K&Z 940187
Now have K&Z 940181 12.08 uV/Wm-2 82.78 Wm-2 / uV
K&Z 940187 9.93 uV/Wm-2 100.70 Wm-2/ uV